a rectangle has a lenght of 4 feet and a perimeter of 14 feet. what is the perimeter of a similar rectangle with a width of 9 feet?
a.36ft
b.42ft
c.108ft
d.126ft

Answer by stanbon(75887)

a rectangle has a length of 4 feet and a perimeter of 14 feet.
Find width:
2(L+W) = 14
L+W = 7
4+W = 7
W = 3
-------------------------------
what is the perimeter of a similar rectangle with a width of 9 feet?
Find the lenght of this rectangle:
L/4 = 9/3
L = 12
------------------------------
Perimeter = 2(L+W) = 2(12+9) = 42 feet
===========================================

a.36ft
b.42ft
c.108ft
d.126ft
===================
cheers,
Stan H.
===================